# The Scene Language: Representing Scenes with Programs, Words, and Embeddings

### v1

@@ Abstract
We introduce the Scene Language, a visual scene representation that concisely and precisely describes the structure, semantics, and identity of visual scenes. It represents a scene with three key components: a program that specifies the hierarchical and relational structure of entities in the scene, words in natural language that summarize the semantic class of each entity, and embeddings that capture the visual identity of each entity. This representation can be inferred from pre-trained language models via a training-free inference technique, given text or image inputs. The resulting scene can be rendered into images using traditional, neural, or hybrid graphics renderers. Together, this forms a robust, automated system for high-quality 3D and 4D scene generation. Compared with existing representations like scene graphs, our proposed Scene Language generates complex scenes with higher fidelity, while explicitly modeling the scene structures to enable precise control and editing. Project page: https://ai.stanford.edu/~yzzhang/projects/scene-language/
@@ 1 Introduction
How do you describe a scene? Imagine that you just traveled to Easter Island and would like to explain to Alice the wondrous scene of Ahu Akivi: “There are seven moai in a row, facing the same direction.” “What is a moai?” Alice asked. “A moai is a stone human figure without legs, but each of them also looks slightly different.” At this point, you realize it seems difficult to precisely explain the scene using natural language alone.
In fact, this example highlights a complete scene representation requires at least three types of complementary information: structural knowledge, which is about the joint distribution of multiple instances, like “seven moai in a row, facing the same direction,” most naturally described as programs; category-level semantics, which may be shared across instances, often described in words, such as “moai”; instance-level intrinsics, tied to the identity of each specific object or part, such as its geometry, color, and texture, which is hard to describe but easy to recognize.
Modern AI techniques provide natural grounding for each of the three modalities, while also falling short of capturing all: in-context learning of pre-trained language models (LMs) enables the inference of domain-specific programs (Brown et al., 2020); LMs capture rich semantic information based on words in natural language; embeddings obtained via techniques like textual inversion (Gal et al., 2023) or low-rank adaptation (Hu et al., 2022) best capture object identity. However, none of these existing representations alone is sufficient for scene generation and editing.
@@
In addition to the representation itself, we propose a training-free inference module using a pre-trained LM as a backbone to infer the Scene Language from texts and images. When provided with a domain-specific language (DSL) for scenes, LMs decompose the task of complex scene generation into simpler tasks of scene component generation by predicting their corresponding modular functions. We also discuss possible neural, traditional, and hybrid graphics engines that render the representation to images. Together, the Scene Language, the inference module, and the renderer lead to a robust system for high-quality, detailed 3D and 4D scene generation and editing.
In summary, our contributions are as follows.
@@
    A training-free method that infers the representation from texts and/or images using pre-trained language models.
@@
    Empirical results on text- and image-conditioned scene generation and editing tasks.
Figure 1: Structured Scene Generation and Editing Using the Scene Language. We develop a scene representation for 3D scene generation and editing tasks. Given textual scene descriptions, the representation can be inferred by a pre-trained large language model, rendered in 3D, and edited following language instructions. The representation contains a program consisting of semantic-aware functions bound to words, providing high interpretability and an intuitive scene-editing interface, and embeddings enabling editing with fine controls, e.g., transferring the style of <z1*> from a user-input image to the generated scene by updating <z1> which controls global attributes of the scene.
@@ 2 Related Work
Visual scene representations are arguably the most fundamental problem in computer vision; thus, for sure, we may not enumerate all related work. As our Scene Language comprises programs, words, and embeddings, we will organize our discussion accordingly into three categories: scene representations that use program-based representations (Section 2.1), semantic graph-based representations (Section 2.2), and a pre-trained generative model’s latent space (Section 2.3).
@@ 2.1 Representing Scenes as Programs
Programs can specify not only the relations among scene components mentioned in Section 2.2, but also structural patterns such as hierarchy and repetitions, making them suitable as explicit descriptions of scene structures. Prior works have proposed to use programs in the form of sequences of execution commands as object-centric representations, followed by neural executors that render the programs into 3D shapes (Tian et al., 2019; Sharma et al., 2018; Deng et al., 2022). In comparison, ShapeAssembly (Jones et al., 2020) introduces higher-level functions with semantically meaningful function names, e.g., “chair” and “back”, to its program representation. Both ShapeAssembly and ours adopt the design principle of function abstraction, which results in clearly stated hierarchy relation among components and better program editability. However, ShapeAssembly uses cuboids as the shape representation and does not model appearance, while ours allows for more precise geometry and appearance modeling using expressive neural embeddings.
All the representations mentioned above require 3D datasets for training. More recently, with the advance of language models (LMs), several methods (Zhou et al., 2024b; Hu et al., 2024; Yamada et al., 2024; Sun et al., 2023; Zhang et al., 2023a; Tam et al., 2024) have proposed to use zero-shot LM inference for generating programs that will be rendered into scenes. These methods operate on top of program syntax from specific graphics renderers such as Blender (Community, 1994), and they do not permit parameters in high-dimensional embedding spaces unlike ours.
@@ 2.2 Representing Scenes with Semantic Graphs
Prior semantic scene representations often adopt a graph to encode semantic scene components, such as objects and parts. In particular, Yuille & Kersten (2006); Huang et al. (2018) propose to employ a parse graph of context-free grammar, using terminal nodes to correspond to objects and their attributes, to represent a scene. Both works employ an analysis-by-synthesis approach to infer the representation from images that heavily rely on domain-specific priors. Alternative representations include scene graph (Johnson et al., 2015; 2018; Gao et al., 2024), where each node in a graph corresponds to an object and an edge corresponds to a pairwise relation, and StructureNet (Mo et al., 2019), which focuses on an object-centric setting and uses nodes for object parts. While these representations preserve the high-level semantics of scenes or objects, they leave out low-level precision; thus, geometric, textural, or relational details that cannot be fully specified by language or hand-crafted rules are often ignored. We address this issue via the inclusion of embeddings.
@@ 2.3 Representing Scenes with Generative Model Latents
The latent space of visual generative models can serve as a representation space for visual scenes. Such latent space can effectively capture the exact visual content of scenes, including geometry and appearance details, and can be either directly inferred, e.g., in variational inference (Kingma, 2014) and model inversion (Zhu et al., 2016). More recently, text-to-image diffusion models have shown remarkable results in image synthesis. This class of models offers several candidate representation spaces including the space of textual embeddings (Gal et al., 2023), low-rank network weights (Hu et al., 2022), full model weights (Ruiz et al., 2023), or noise vectors in the diffusion process (Song et al., 2021; Mokady et al., 2023; Ho et al., 2020). However, such representations typically do not offer interpretable semantics or explicitly encode hierarchical scene structures unlike ours.
Figure 2: Scene Language Overview. A Scene Language represents a scene with three components: a program consisting of entity functions, a set of words (e.g., pawn) denoting the semantic class of the entity functions, and a list of embeddings (e.g., <z1>) capturing the identity of each entity in the scene. Each entity function is bound with an entity class name given by a word, and maps an input embedding to an output entity of that class. Executing the program effectively computes all entities; the computation graph is shown on the right. Entity dependency, as indicated by arrows, reflects the hierarchical relation of entities in a scene. See Section 3.1 for representation definitions and Section 3.2 for program syntax. The program shown is converted from our inference method output, with text prompt “a chessboard at game start”; raw outputs in Section F.1.
Section 3.1 Section 3.2 Definition
@@
$f_{w}:z,\gamma\mapsto h$ <entity-func> Entity function mapping embeddings to entity
@@
$w$ Word Word describing semantics
@@
$z$ Embedding Embedding specifying entity identity
$\gamma$ List[Embedding ] Embeddings specifying identities of descendent entities
@@
Table 1: Summary of Notations in Sections 3.1 and 3.2.

### v2

@@ Abstract
We introduce the Scene Language, a visual scene representation that concisely and precisely describes the structure, semantics, and identity of visual scenes. It represents a scene with three key components: a program that specifies the hierarchical and relational structure of entities in the scene, words in natural language that summarize the semantic class of each entity, and embeddings that capture the visual identity of each entity. This representation can be inferred from pre-trained language models via a training-free inference technique, given text or image inputs. The resulting scene can be rendered into images using traditional, neural, or hybrid graphics renderers. Together, this forms an automated system for high-quality 3D and 4D scene generation. Compared with existing representations like scene graphs, our proposed Scene Language generates complex scenes with higher fidelity, while explicitly modeling the scene structures to enable precise control and editing.
Figure 1: Structured Scene Generation and Editing Using the Scene Language. We develop a scene representation for 3D scene generation and editing tasks. Given textual scene descriptions, the representation can be inferred by a pre-trained large language model, rendered in 3D, and edited following language instructions. The representation contains a program consisting of semantic-aware functions bound to words, providing high interpretability and an intuitive scene-editing interface, and embeddings enabling editing with fine controls, e.g., transferring the style of <z1*> from a user’s image to the output by updating <z1> which controls global attributes of the scene. Project page: https://ai.stanford.edu/~yzzhang/projects/scene-language/.
Figure 2: Overview. A Scene Language represents a scene with three components: a program consisting of entity functions, a set of words (e.g., pawn) denoting the semantic class of the entity functions, and a list of embeddings (e.g., <z1>) capturing the identity of each entity in the scene. Each entity function is bound with an entity class name given by a word, and maps an input embedding to an output entity of that class. Executing the program evaluates entity functions to compute the full set of entities in the scene. The resulting computation graph, illustrated on the right, captures the dependency structure among entities (indicated by arrows). The program shown is converted from our text-conditioned inference method’s output, with details included in Sec. F.2.1; it is written in Lisp-style syntax for brevity and is implemented with Python in practice (Sec. 3.2).
@@ 1 Introduction
How do you describe a scene? Imagine that you just traveled to Easter Island and would like to explain to Alice the wondrous scene of Ahu Akivi: “There are seven moai in a row, facing the same direction.” “What is a moai?” Alice asked. “A moai is a stone human figure without legs, but each also looks slightly different." At this point, you realize it seems difficult to precisely explain the scene using natural language alone.
In fact, this example highlights a complete scene representation requires at least three types of complementary information: structural knowledge, which is about the joint distribution of multiple instances, like “seven moai in a row, facing the same direction,” most naturally described as programs; category-level semantics, which may be shared across instances, often described in words, such as “moai”; and instance-level intrinsics, tied to the identity of each specific object or part, such as its geometry, color, and texture, which is hard to describe but easy to recognize.
Modern AI techniques provide natural grounding for each of the three modalities, while also falling short of capturing all: in-context learning of pre-trained language models (LMs) enables the inference of domain-specific programs [3]; LMs themselves capture rich semantic information based on words in natural language; and embeddings obtained via techniques like textual inversion [6] or low-rank adaptation [9] best capture object identity. However, none of these existing representations alone is sufficient for scene generation and editing.
@@
In addition to the representation itself, we propose a training-free inference module using a pre-trained LM as a backbone to infer the Scene Language from texts and images. When provided with a domain-specific language (DSL) for scenes, LMs decompose the task of complex scene generation into simpler tasks of scene component generation by predicting their corresponding modular functions. We also discuss possible neural, traditional, and hybrid graphics engines that render images based on the representation. Together, the Scene Language, the inference module, and the renderer lead to a system for high-quality, detailed 3D and 4D scene generation and editing.
In summary, our contributions include
@@
    A training-free method that infers the representation from texts and/or images using pre-trained LMs.
@@
    Empirical results on text- and image-prompted scene generation and editing tasks.
@@ 2 Related Work
Visual scene representations are arguably the most fundamental problem in computer vision; thus, for sure, we may not enumerate all related work. As our Scene Language comprises programs, words, and embeddings, we organize our discussion accordingly into three categories: scene representations using program-based representations (Sec. 2.1), semantic graph-based representations (Sec. 2.2), and a pre-trained generative model’s latent space (Sec. 2.3).
@@ 2.1 Representing Scenes as Programs
Programs can specify not only the relations among scene components mentioned in Sec. 2.2, but also structural patterns such as hierarchy and repetitions, making them suitable as explicit descriptions of scene structures. Prior works have proposed to use programs in the form of sequences of execution commands as object-centric representations, followed by neural executors that render the programs into 3D shapes [36, 30, 5]. In comparison, ShapeAssembly [17] introduces higher-level functions with semantically meaningful function names, e.g., “chair” and “back”, to its program representation. Both ShapeAssembly and ours adopt the design principle of function abstraction, which results in clearly stated hierarchy relation among components and better program editability. However, ShapeAssembly uses cuboids as the shape representation and does not model appearance, while ours allows for more precise geometry and appearance modeling using expressive neural embeddings.
All representations above require 3D datasets for training. More recently, with the advance of language models (LMs), several methods [43, 10, 37, 33, 39, 34] have proposed to use zero-shot LM inference for generating programs that will be rendered into scenes. These methods operate on top of program syntax from specific graphics renderers such as Blender [4], and they do not permit parameters in high-dimensional embedding spaces unlike ours.
@@ 2.2 Representing Scenes with Semantic Graphs
Prior semantic scene representations often adopt a graph to encode semantic scene components, such as objects and parts. In particular, [38, 11] propose to employ a parse graph of context-free grammar, using terminal nodes to correspond to objects and their attributes, to represent a scene. Both works employ an analysis-by-synthesis approach to infer the representation from images that heavily rely on domain-specific priors. Alternative representations include scene graph [15, 16, 7], where each node in a graph corresponds to an object and an edge corresponds to a pairwise relation, and StructureNet [22], which focuses on an object-centric setting and uses nodes for object parts. While these representations preserve the high-level semantics of scenes or objects, they leave out low-level precision; thus, geometric, textural, or relational details that cannot be fully specified by language or hand-crafted rules are often ignored. We address this issue via the inclusion of embeddings.
@@ 2.3 Representing Scenes with Generative Model Latents
The latent space of visual generative models can serve as a representation space for visual scenes. Such latent space can effectively capture the exact visual content of scenes, including geometry and appearance details, and can be either directly inferred, e.g., in variational inference [19] and model inversion [45]. More recently, text-to-image diffusion models have shown remarkable results in image synthesis. This class of models offers several candidate representation spaces including the space of textual embeddings [6], low-rank network weights [9], full model weights [29], or noise vectors in the diffusion process [32, 24, 8]. However, such representations typically do not offer interpretable semantics or explicitly encode hierarchical scene structures unlike ours.
3 The Scene Language
We aim to design a visual scene representation that encodes the structure, semantics, and visual content of scenes. Towards this goal, we propose the Scene Language, which represents a scene with three components: a program that encodes scene structure by specifying the existence and relations of scene components, which we will refer to as entities; words in natural language that denote the semantic group of each entity in the scene; and neural embeddings that pertain the low-level visual details and identities of the entities by permitting an expressive input parameter space. In the following, we will first give a formal definition of the representation (Sec. 3.1), and then introduce a domain-specific language (DSL) (Sec. 3.2) as its realization.
Sec. 3.1 Sec. 3.2 Definition
@@ 3 The Scene Language
$f_{w}:z,\gamma\mapsto h$ <entity-func> Mapping embeddings to entity
@@
$w$ Word Describing semantics
@@
$z$ Embedding Specifying entity identity
$\gamma$ List[Embedding ] Specifying descendent entities’ identities
@@
Table 1: Summary of Notations in Secs. 3.1 and 3.2.
Renderer Examples
Primitive-based Light transport Shape and BRDF LM inference
Asset-based Ray tracing Asset metadata LM inference
SDS-based Gaussian splatting Gaussian parameters SDS optimization
T2I model Feed-forward pass Text embeddings in $\mathcal{Z}$ CLIP encoding
Table 2: Graphics Renderers Examples.
3.1 Formal Definition
$\Phi(s):=\left(W,P,Z\right).$ (1)
The following expands on how functions from $P$ are defined and the program execution procedure to compute the represented scene $s$. Notations are summarized in Tab. 1.
Entity Function Definitions.
where $h^{(i)}=f_{w^{(i)}}(z^{(i)},\gamma^{(i)})$ is a sub-entity of $h$ and $t^{(i)}(z)$ specifies its pose.
Program Execution.
3.2 Scene Language as Programming Language
We now concretize the definition in Sec. 3.1 with a domain-specific language (DSL) specified in Sec. B.1. We introduce macro operations union for $\Psi_{\text{union}}$, union-loop which calls union on entities evaluated in a for-loop, and transform for $\Psi_{\text{transform}}$. Together with call which applies function calls, these four macro operations fully define entity functions in the DSL. Entity functions are identified with the associated words in the DSL via two special forms: bind, which binds entity function $f_{w}$ to word $w$, and retrieve, which retrieves $f_{w}$ given $w$ and applies $f_{w}$ on actual embedding parameters. If $w$ is never bound to a function, it corresponds to an entity function with no sub-entities (i.e., $N=0$ in Eq. 2), in which case call returns a primitive entity with no children.
The data type of an entity $h$ (Eq. 2) is denoted as Entity. It stores two data fields Word and Embedding, describing the entity’s semantic group and identity respectively, and stores each children entity together with its pose in the frame of $h$. In particular, Embedding captures the visual details and requires a highly expressive representation, e.g., neural embeddings. We employ the textual embedding space of OpenCLIP-ViT/H [13] for parameterization, denoted as $\mathcal{Z}_{\text{CLIP}}$. It offers the advantage that embeddings can be either encoded directly from natural language or inferred from images with Textual Inversion [6]. Tab. 1 summarizes the expressions and data types in accordance with the notations from Sec. 3.1.
4 Rendering
Renderer Specifications.
Renderer Instantiations.
An example renderer instantiation is with Score Distillation Sampling (SDS) [25] guidance, where $\Theta$ is a differentiable 3D representation, and we specify $g_{\text{reparam}}$ as follows. Recall that an entity is associated with a Word value, e.g., moai for one statue from Figs. 1 and 3, and an Embedding value, e.g., <z2>. For each primitive entity (i.e., an entity with no children), given these two value fields together with the fields of its ancestors, we use a manually specified language template $c$, or a <z2> moai, in the style of <z1>, 3D model in this example, to embed them into $z=g_{\text{CLIP}}(c)\in\mathcal{Z}_{\text{CLIP}}$ where $g_{\text{CLIP}}$ is the pre-trained CLIP text encoder. Then, $g_{\text{reparam}}:\mathcal{Z}_{\text{CLIP}}\rightarrow\Theta$ corresponds to the SDS-guided optimization to find a solution in $\Theta$ that aligns with the input condition $z\in\mathcal{Z}_{\text{CLIP}}$. Output 3D scenes can be personalized by editing embeddings, e.g., <z1> which controls the global style in Fig. 1.
For the underlying 3D representation, we use 3D Gaussian Splatting [18] where images are rendered by splatting a set of 3D Gaussians onto the image plane; other differentiable 3D representations such as neural fields will also be suitable. We base our implementation on GALA3D [44], and use MVDream [31] and a depth-conditioned ControlNet [40] for guidance.
We refer to the renderer above as the Gaussians renderer. Other possible renderers include primitive-based renderers, such as Mitsuba [14] with graphics primitives of cubes, spheres, and cylinders; asset-based game engines, such as MineCraft [23]; and feed-forward inference of layout-conditioned text-to-image (T2I) diffusion models, such as MIGC [42], which achieves 2D bounding box conditioning by controlling attention layers from Stable Diffusion [28]). Tab. 2 shows a summary; details are in Sec. B.2.
5 Inference via Pre-Trained Language Models
LMs have shown remarkable capability in code generation with common programming languages such as Python. In our implementation, we prompt LMs to generate Python scripts. We prompt the LM with the input condition, i.e., a scene description in texts or an image; a Python script of helper functions converted from the DSL in Sec. 3.2; and an example script using the helper functions. We use Claude 3.5 Sonnet [1] for all experiments for our method and LM-dependent baselines. Full LM prompts are in Sec. F.1.
Function arguments in the LM-generated scripts, which are numeric values or string tokens, are converted to embeddings from $\mathcal{Z}_{\text{CLIP}}$ (Sec. 3.2) using language templates and the CLIP text encoder $g_{\text{CLIP}}$. For example, in the raw LM output, function calls for white pieces in Fig. 2 have input attribute {"color":(.9,.9,.9)}, and we prompt LM to describe the color value as a word, and feed the word into $g_{\text{CLIP}}$ to compute <z68>. For image-conditioned tasks, for each primitive entity in the execution output of $P$, we first use GroundingSAM [20, 27] to segment out the region defined by the word associated with the entity. We then use Textual Inversion [6] to optimize an embedding to reconstruct the cropped image with the diffusion model training objective. The full process is deferred to Sec. C.2.
Figure 4: Text-Prompted Scene Generation. Ours produce more accurate outputs than using no intermediate representation (MVDream) or scene graph (GraphDreamer).
Numeric Prompts Generic Prompts
Methods Alignment ($\uparrow$) CLIP ($\uparrow$) Count ($\uparrow$) Alignment ($\uparrow$) CLIP ($\uparrow$)
GraphDreamer [7] $3.56$±$\pm 7.38$ $0.297$±$\pm 0.085$ 0.11 $15.75$±$\pm 20.39$ $0.230$±$\pm 0.097$
MVDream [31] $10.79$±$\pm 12.83$ $0.312$±$\pm 0.046$ 0.11 $35.13$±$\pm 31.90$ $\mathbf{0.328}$±$\pm\mathbf{0.031}$
Ours $\mathbf{85.65}$±$\pm\textbf{13.71}$ $\mathbf{0.351}$±$\pm\textbf{0.051}$ $\mathbf{1.0}$ $\mathbf{52.13}$±$\pm\mathbf{26.83}$ $\mathbf{0.328}$±$\pm\mathbf{0.030}$
Table 3: Evaluation for text-prompted 3D generation. We report the following metrics averaged across $9$ numeric and $8$ generic prompts: the percentages of user preferences for prompt alignment, CLIP similarity scores, and counting accuracy (numeric prompts only, 00 for inaccurate and $1$ for accurate). Here $\pm$ denotes standard deviation across scenes.
Figure 5: Renderings Across Graphics Renderers. Different renderers produce renderings that adhere to the same representation and are therefore visually aligned, while each exhibits a different imaging style. Text inputs are shown at the subfigure bottoms.
Figure 6: Text-Prompted 4D Scene Generation. The proposed representation may also capture scene dynamics and be applied for synthesizing 4D scenes. It explicitly represents the temporal correspondence of an entity in a dynamic scene. Each colored trajectory denotes tracking of a temporally moving point. Compared to 4D-fy, our method produces more salient motion.
Figure 7: Scene Editing with Language Instructions. Our representation is editing-friendly. Bottoms of each row show initial scene descriptions and editing instructions in the format of user text prompts. We prompt an LM to infer the initial Scene Language as well as the edits (shown with texts in red), and show image renderings with two renderers.
Figure 8: Scene Editing with Image Instructions. This figure shows style transfer results under the same setup as in Fig. 1. Replacing embeddings of target entities in a Scene Language with a new embedding specified by a user image achieves the editing effect.
Figure 9: Image-Prompted Scene Generation. Compared to scene graphs from GraphDreamer, programs from our representation encode additional scene structure, e.g., repetitions, and specify geometric entity relations more precisely. Embeddings from ours further enable visual identity preservation in the outputs.
6 Experiments
6.1 Text-Prompted Generation and Editing
Baselines.
To evaluate our representation in text-prompted 3D generation task, we compare our inference pipeline with 3D scene generation methods using alternative intermediate representations, e.g., scene graph. In particular, we compare with GraphDreamer [7] as an exemplar approach, which generates scene graphs from texts via LM prompting and then synthesizes scenes conditioned on the graphs via SDS guidance. We further ablate the role of structural representation in this task by comparing ours with the backbone of our SDS-based renderer, MVDream [31], as a direct scene generation approach. Details are in Sec. C.3.
Methods CLIP [13] ($\uparrow$) Dynamic [12] ($\uparrow$)
4D-fy [2] 0.352 $0.2\%$
Ours 0.341 5.9%
Methods LPIPS [41] ($\downarrow$)
GraphDreamer [7] 0.811
Ours 0.681
Table 4: Evaluation. Left: For text-prompted 4D generation, 4D-fy achieves higher prompt alignment measured with CLIP similarity but has small motion. Right: For image-prompted scene generation, our results are better aligned with input image prompts compared to GraphDreamer.
Metrics.
We conduct a user study where $103$ users are asked to choose one of the three videos, showing 360-renderings from ours and two baselines in a randomized order, that aligns the best with text prompts. Following [7], for each scene, we compute the cosine similarity of CLIP [26, 13] embeddings between a rendering and the text prompt and report the average over $120$ viewpoints. We further report whether outputs have the correct object counts. Details are deferred to Sec. C.4.
Results.
Tab. 3 shows that our method achieves a more favorable prompt alignment than the baselines as measured in the user study and CLIP similarity, and has a clear advantage in counting accuracy.
Fig. 4 and Sec. D.1 show qualitative results of ours using the SDS-based renderer. Compared to the direct 3D scene generation method MVDream, our approach is compositional and adheres more closely to input prompts in scenes involving multiple objects. Compared to a scene graph representation, where entity relations are restricted to be between two objects and are bottlenecked by the coarseness of natural language descriptions, e.g., “aligned in a row”, a program-based representation offers more flexible and precise specifications for relations, e.g., the particular coke can arrangement in Fig. 4. This brings the practical benefit of offloading the burden of generating scenes involving complex entity relations from the T2I model (used for SDS guidance in both ours and GraphDreamer) towards LM, leading to accurate and detailed generation results.
Across Graphics Renderers.
Fig. 5 shows that our representation can be rendered with different renderers (Sec. 4), producing aligned outputs in different visual domains.
4D Generation.
The inference method from Sec. 5 can be applied to 4D scene generation. The 4D scene representation is identical to Eq. 1, except that there is an additional 4D entity function in the program $P$.
We compare with 4D-fy [2] as an exemplar text-to-4D generation method. We measure the prompt alignment of output dynamic renderings using CLIP similarity between rendered frames and input texts, and the motion magnitude using dynamic degrees [12] computed using RAFT [35] optical flow estimation over neighboring frames. 4D-fy results have a higher per-frame alignment score than ours but present small dynamic motions. See Fig. 6, the webpage, and Sec. D.2 for further results.
Scene Editing.
Scenes synthesized from our proposed representation can further be edited by prompting LM with its previously generated script and an editing instruction in natural language instruction. Results are shown in Fig. 7. Our representation provides an interpretable and intuitive interface for scene editing, as functions have explicit semantic meanings associated with words, and function reuse significantly improves program readability. Furthermore, since the structure of programs reflects the structure of scenes, editing program parameters leads to precise changes in the scenes while preserving the original structure, e.g., the circular arrangement of staircases in Fig. 7. Desirable editing effects involving multiple primitives, or all staircases in this example, can be effectively achieved via only minor changes in the program space. Finally, the program structure itself, such as the function header in the Jenga set example, can be adjusted for editing to achieve localized edits that only affect relevant parts of the scene. Further results are shown in Sec. D.3.
6.2 Image-Prompted Generation and Editing
Texts provide an easy user interface for scene generation as shown in Sec. 6.1, and images allow users to specify further intentions that could be hard to describe with natural language alone. Integrating neural embeddings increases the expressivity of the Scene Language and allows it to encode visual details from image prompts, and as a result, enhances our method’s controllability in image-prompted scene generation and editing tasks.
Generation.
Fig. 9 shows results on the task of generating 3D scenes consistent with the structure and content from a user-provided image, termed as the “inverse semantics” task in GraphDreamer [7]. We report the LPIPS [41] score, i.e., the AlexNet [21] feature dissimilarity, between an input image and an output rendering, averaged over 3 scenes with $20$ viewpoints each in Tab. 4. Embeddings allow our representation to encode visual details from inputs, such as the Coke can texture in Fig. 9, which is not captured in scene graphs from GraphDreamer using per-object natural language descriptions. This is reflected quantitatively in Tab. 4.
Editing.
Our representation applies to image-prompted editing tasks such as style transfer (Fig. 1). Edits can target partial or full scenes as shown in Fig. 8(a) and (b), respectively, by updating embeddings for corresponding entities.
7 Conclusion
We have introduced a visual scene representation, termed the Scene Language, which encodes three key aspects of visual scenes: scene structure, such as hierarchy and repetition, specified via programs; semantics of individual components succinctly summarized via words; and identities of each component precisely captured via neural embeddings. We formalize the representation as a programming language defined using a DSL. We show that the Scene Language can be efficiently inferred from both text and image inputs leveraging pre-trained language models. Once the program is executed, the resulting scene can be rendered into images using a variety of graphics renderers. Our Scene Language produces 3D and 4D scenes with high fidelity and enables easy and precise editing.
Acknowledgments.
This work is in part supported by AFOSR YIP FA9550-23-1-0127, ONR N00014-23-1-2355, ONR YIP N00014-24-1-2117, ONR MURI N00014-22-1-2740, and NSF RI #2211258 and #2338203. YZ is in part supported by the Stanford Interdisciplinary Graduate Fellowship.
Appendix A Overview
This file contains representation details (Appendix B), experiment details (Appendix C), extended experiment results (Appendix D), discussions of limitations (Appendix E), and full language model prompts and outputs (Appendix F).
Appendix B Representation Details
B.1 Domain-Specific Language
The complete DSL is listed in Tab. 5. We explain the four macros introduced in Sec. 3.2, also listed in Tab. 5, as follows.
  • Macro call retrieves <entity-func> bound to the input word, applies the function on the input embeddings, and outputs a data object of type Entity. Specifically, (car embedding-list) is the embedding of the output entity corresponding to $z$ from Eq. 2, and (cdr embedding-list) is the embeddings of its descendent entities corresponding to $\gamma$.
  • Macro union-loop applies union in a for loop.
  • Macro transform pairs an entity with its pose.